OpenStudy (anonymous):

Paul needs 2 1/4 yards of fabric to make a tablecloth. How many tablecloths can he make from 18 1/2 yards of fabric? A.72 B.10 C.9 D.8

OpenStudy (anonymous):

5

OpenStudy (anonymous):

nvm its c

OpenStudy (anonymous):

hi again diana.

OpenStudy (anonymous):

long time no see :)

OpenStudy (anonymous):

lol hey hsmt! help again ? i hate this worksheet!

OpenStudy (anonymous):

lol.. some worksheets are killers arent they? sure i can help you

OpenStudy (anonymous):

Thankkkk youuuuu!!

OpenStudy (anonymous):

seems like hundreds of aimless problems

OpenStudy (anonymous):

sure. lets make this a little easier to start.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

if i had 10 yards of fabric and i needed two yards for my tablecloth, how many tablecloths could i make from the 10 yards. :)

OpenStudy (anonymous):

about 3 so it would be 2 because no one wants 3.3 of a table cloth

OpenStudy (anonymous):

i think lol

OpenStudy (anonymous):

you have 10 yards

OpenStudy (anonymous):

lets say you have 10 soda pops and you have 2 people and you want to DIVIDE the soda pops equally among the 2 people. How many sodas would each person get?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

i feel so dumb -.-

OpenStudy (anonymous):

yes.. perfect. note that it is the total number of sodas divided by the number of people.. you are doing fine. :)

OpenStudy (anonymous):

now, you have a total of 18 1/2 yds of fabric and it takes 2 1/4 yards per tablecloth. what operation would i use here (addition, multiplication, division)?

OpenStudy (anonymous):

or you have 18 1/2 soda pops and 2 1/4 people

OpenStudy (anonymous):

i think divison and i did what you told me on the last thing and i got like 8. something so i thought it might be 9 or 8

OpenStudy (anonymous):

show me what you did to get 8

OpenStudy (anonymous):

waiiiittt do you have divide with 18 and 2 so it would be 9 and something

OpenStudy (anonymous):

close. what you do is turn the two mixed numbers into improper fractions. if you need help with this tell me and i will show you how.

OpenStudy (anonymous):

yeah i think thats what i did and i came up with 9/4 37/2

OpenStudy (anonymous):

yes.. very nice :)

OpenStudy (anonymous):

then what did you do?

OpenStudy (anonymous):

i fliped the 37/2 with 2/37 i crossed out the 2 and for so i have 9/2 1/37

OpenStudy (anonymous):

now, the question here is which one is the dividing fraction? it is the total, divided by the amount needed to make 1 tablecloth

OpenStudy (anonymous):

you are close... very close

OpenStudy (anonymous):

Good and thats where i got lost i dont know what to do anymore do i divde do i multiply?

OpenStudy (anonymous):

ok, the total available material is 18 1/4 like the total available soda pops was 10.

OpenStudy (anonymous):

to get the number of table cloths, you divide the total available fabric by the number of yards for one table cloth. just like you divided the 10 soda pops amongst 2 people.

OpenStudy (anonymous):

so the dividing fraction is 9/2.

OpenStudy (anonymous):

so 9/2 divided by 1/37?

OpenStudy (anonymous):

not quite...

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

i figured that wouldnt work so what next?

OpenStudy (anonymous):

try 37/2 divided by 9/4

OpenStudy (anonymous):

ohhh yeaaahhh thats how i got 8 2/9 thats why i said i got that but decide to round to 9

OpenStudy (anonymous):

the total available material divided by the amount of material required for EACH tablecloth. if you are in doubt ask question s. lots of them if necessary...:)

OpenStudy (anonymous):

so i couldnt decide!

OpenStudy (anonymous):

so its 9?

OpenStudy (anonymous):

right. your worksheet is a little mean there.

OpenStudy (anonymous):

lol thank you soo much!!

OpenStudy (anonymous):

well, if you round 8 2/9

OpenStudy (anonymous):

so wait is it 8 or 9

OpenStudy (anonymous):

it would round down to 8 since 2/9 is less than 1/2

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

you are so welcome. while i am here, do you have any other questions?

OpenStudy (anonymous):

well i ahve the question that i have been wondering for really long but i feel too dumb to ask but how is 2/9 smaller then 1/2 and 1/4 is smaller too like i dont get the fractions

OpenStudy (anonymous):

ok... good question.

OpenStudy (anonymous):

the answer is in equivalent fractions. for instance 1/2=2/4=3/6=4/8=9/18. hmmm... why 9/18 you say... glad you asked

OpenStudy (anonymous):

2/9=4/18

OpenStudy (anonymous):

i got the above equivalent fraction (same value- different numerator and denominator) by multiplying the numerator and denominator of 2/9 by 2. if i multiplied 2/9 by 3/3, then i would get an equivalent fraction of 6/27. does that make sense?

OpenStudy (anonymous):

notice that to change 1/2 to 2/4, you can do so by multiplying 1/2 by the fraction 2/2 and multiplying numerators and denominators straight across

OpenStudy (anonymous):

kinda i also have another question Sasha finished mowing lawns at 4:15 PM. It took 1 3/4 h to mow the first lawn. The second and third lawns she mowed each took 1 1/5 h. She took one 45-min break. When did Sasha begin mowing the first lawn? D:

OpenStudy (anonymous):

2/2=1, so i didnt change the value of the fraction.

OpenStudy (anonymous):

nevermind i firgured it out the questtion not the fractions

OpenStudy (anonymous):

let what i have said just kind of mull around in your mind. you will figure this out. for instance, the tablecloth question is asking how many 2 1/4's are there in 18 1/2?

OpenStudy (anonymous):

division by fractions is subtle and strangely annoying... lol

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

like if you ask how many 1/2's are there in 3, it would be 3/1 divided by 1/2=6. so there are six 1/2's in 3 or six 50 cent pieces in $3.00

OpenStudy (anonymous):

you use fractions all the time when you deal with money.. that is the frustrating part. you actually know fractions very well.

OpenStudy (anonymous):

this is so confussing

OpenStudy (anonymous):

i know. the rules i have given you will work for all fractions in division. sometimes word problems make it even more confusing.

OpenStudy (anonymous):

tell me what is confusing you at this moment and i will try to clarify for u. there are no dumb questions except the ones that need answers and remain unasked.

OpenStudy (anonymous):

im sorry but everything you just said......

OpenStudy (anonymous):

nevermind i reread it and its kinda clear now

OpenStudy (anonymous):

ok.. let me ask u something

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

when you take tests, do you get really anxious?

OpenStudy (anonymous):

and then it gets hard to think?

OpenStudy (anonymous):

not really i do when i hit the start button but once i read all the questions i am like okay i can do this and yeah i have to think it through but this was the hardest one.

OpenStudy (anonymous):

ok... good... glad to hear it. if you think like that, it prevents adrenaline from rushing into your body and clouding your ability to think

OpenStudy (anonymous):

is there anything else i can help with?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

thanks alot for wasting your time on me (:

OpenStudy (anonymous):

you are an sweet intelligent being. there was no waste of time here. just because everything didnt click this time, does not mean it will not click later on.. just give it some time. the mind is a funny thing. :)

OpenStudy (anonymous):

it absorbs info at its own rate and in its own time. you'll get there.

OpenStudy (anonymous):

:D Thanks

OpenStudy (anonymous):

thats sweet lol

OpenStudy (anonymous):

my pleasure..

OpenStudy (anonymous):

you take care

OpenStudy (anonymous):

i hope to run into you again some time.